what mass of FeCl3 is required to prepare 100mL of 0.250 M FeCl3 solution?
1
Expert's answer
2013-05-17T10:41:10-0400
To calculate the amount of FeCl3, we can use the formula:
Cm = n(FeCl3)/Vsolution
Where
Cm - molar concentration
n - amount in moles
V - volume of solution
n(FeCl3) = Cm * Vsolution
n(FeCl3) = 0.250 M * 0.1L = 0.025 moles
To calculate the mass of FeCl3, we can use the formula:
m(FeCl3) = n(FeCl3) * M(FeCl3)
Where
M - molar mass
n - amount in moles
M(FeCl3) = Ar(Fe) +3*Ar(Cl)
M(FeCl3) = 56 + 3 * 35.5 = 162.5 g/moles
m(FeCl3) = 0.025 moles * 162.5 g/moles = 4.0625 g

So, the final answer is 4.0625 g of FeCl3